Question 1Coin and die
A fair coin is tossed and a fair 4-sided die, numbered 1 to 4, is thrown. Find the probability of getting a head and a 3.
Show the worked solution
- Step 1: Find each probability, then combine them: both must happenP(head)=21P(3)=41(P=21×41P=81
Question 2Coin and die
A fair coin is tossed and a fair 8-sided die, numbered 1 to 8, is thrown. Find the probability of getting a tail and a 5.
Show the worked solution
- Step 1: Find each probability, then combine them: both must happenP(tail)=21P(5)=81(P=21×81P=161
Question 3Both, one but not the other, neither
The probability that it rains tomorrow is 0.2 and the probability that the bus is late is 0.9. The two events are independent. Find the probability that it rains tomorrow but the bus is not late.
Show the worked solution
- Step 1: P(not happening) = 1 − P(happening)P(second does not occur)=1−0.9P(second does not occur)=0.1
- Step 2: Combine the two probabilities: both must happen(P=0.2×0.1P=0.02
Question 4Both, one but not the other, neither
The probability that Ali passes his driving test is 0.9 and the probability that Ben passes his driving test is 0.15. The two events are independent. Find the probability that neither event occurs.
Show the worked solution
- Step 1: P(not happening) = 1 − P(happening)P(first does not occur)=1−0.9P(first does not occur)=0.1P(second does not occur)=1−0.15P(second does not occur)=0.85
- Step 2: Combine the two probabilities: both must happen(P=0.1×0.85P=0.085
Question 5At least one
The probability that Ali passes his driving test is 0.4 and the probability that Ben passes his driving test is 0.5. The two events are independent. Find the probability that at least one of the two events occurs.
Show the worked solution
- Step 1: The probability that each event does not occurP(first does not occur)=1−0.4P(first does not occur)=0.6P(second does not occur)=1−0.5P(second does not occur)=0.5
- Step 2: Neither occurs: combine the two(P(neither)=0.6×0.5P(neither)=0.3
- Step 3: At least one occurs: use the result for neither(P=1−0.3P=0.7
Question 6At least one
The probability that Ali passes his driving test is 0.25 and the probability that Ben passes his driving test is 0.4. The two events are independent. Find the probability that at least one of the two events occurs.
Show the worked solution
- Step 1: The probability that each event does not occurP(first does not occur)=1−0.25P(first does not occur)=0.75P(second does not occur)=1−0.4P(second does not occur)=0.6
- Step 2: Neither occurs: combine the two(P(neither)=0.75×0.6P(neither)=0.45
- Step 3: At least one occurs: use the result for neither(P=1−0.45P=0.55